OpenStudy (anonymous):

"If x>2, then x^2-x-6/x^2-4=??" (Please show your work, Thanks!)

jimthompson5910 (jim_thompson5910):

what do you get when you factor the numerator?

OpenStudy (anonymous):

Umm...hehe...Don't the two x^2's cancel each other out...? (sorry... :/ )

jimthompson5910 (jim_thompson5910):

no that's a common misconception, but they don't

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

what two numbers multiply to -6 and add to -1?

jimthompson5910 (jim_thompson5910):

if you're not sure, list out the possible ways to factor -6 -6 = -6*1 -6 = -3*2 -6 = -2*3 -6 = -1*6 now which pair of factors add to -1 ?

OpenStudy (anonymous):

-2*3

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

-2+3 = 1 not -1

OpenStudy (anonymous):

OH the other one: -3*2 but why -6 and -1?

jimthompson5910 (jim_thompson5910):

I was just listing all the possible ways to multiply two whole numbers to get -6

jimthompson5910 (jim_thompson5910):

yes -3+2 = -1

jimthompson5910 (jim_thompson5910):

the two numbers are -3 and 2 so x^2 - x - 6 factors to (x-3)(x+2)

jimthompson5910 (jim_thompson5910):

now do the same for x^2-4 and think of x^2-4 as x^2+0x-4

OpenStudy (anonymous):

For Numerator u can always use this formula for solving such equations: ax^2+bx+c=0 then roots are x=[-b+Squrt(b^2-4ac)]/2a & [-b-sqrt(b^2-4ac)]/2a For Denominator There is an identity x^2-a^2=(x+a)(x-a) Then cancel out the common terms This should help

jimthompson5910 (jim_thompson5910):

so what do you get when you factor x^2-4 or x^2+0x-4

OpenStudy (anonymous):

So...now find a factor for -4?

jimthompson5910 (jim_thompson5910):

correct list all the possible ways to multiply to -4

jimthompson5910 (jim_thompson5910):

see which factorization pair adds to 0

OpenStudy (anonymous):

Proceed @jim_thompson5910 is saying right now but after completing this sum . Remember those identities . it will be helpful for future problem solving

OpenStudy (anonymous):

2*-2= -4 and 2+(-2)=0

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so x^2 - 4 factors to (x+2)(x-2)

jimthompson5910 (jim_thompson5910):

you now have (x-3)(x+2) ---------- (x+2)(x-2)

jimthompson5910 (jim_thompson5910):

I'm sure you see what cancels

OpenStudy (anonymous):

(x+2)/(x-2) cancels

OpenStudy (anonymous):

so now left with (x-3)/(x+2)?

jimthompson5910 (jim_thompson5910):

the x+2 terms cancel I think you meant to say

jimthompson5910 (jim_thompson5910):

only identical terms can cancel because x/x = 1

OpenStudy (anonymous):

right, sorry

jimthompson5910 (jim_thompson5910):

so you're left with x-3 ------- x-2

OpenStudy (anonymous):

Kay

OpenStudy (anonymous):

Thanks a lot @jim_thompson5910 ! Really appreciate it!

OpenStudy (anonymous):

Now u can have a look at these :)

jimthompson5910 (jim_thompson5910):

you're welcome